OpenStudy (anonymous):

Suppose y varies directly with x. If x = -8 when y = 1, find the value of y when x = 2. Select one: a. -3/5 b. -4 c. 2 d. -1/4

OpenStudy (cj49):

y=kx 1=k-8 k=1/-8

OpenStudy (anonymous):

ok

OpenStudy (cj49):

y=1/-8(2) y=-1/4

OpenStudy (anonymous):

So that's the formula right?

OpenStudy (anonymous):

Thanks.

OpenStudy (cj49):

np....yes

OpenStudy (anonymous):

Suppose y varies directly with x. If x = 7 when y = -2, find the value of x when y = -6. y=kx -2 = k + 7 k = -2/7 y = -2/7(-6) y = 1.7

OpenStudy (cj49):

y=kx is the formula

OpenStudy (anonymous):

?? that's not right..

OpenStudy (anonymous):

what i did ^^

OpenStudy (cj49):

its correct !!

OpenStudy (anonymous):

for real? well my options are this: a. -3 b. 12 c. 21 d. 7

OpenStudy (cj49):

@aryandecoolest

OpenStudy (cj49):

idk wat ya doing wrong here

OpenStudy (anonymous):

in which question among those u need help?

OpenStudy (cj49):

2nd one

OpenStudy (anonymous):

Suppose y varies directly with x. If x = 7 when y = -2, find the value of x when y = -6.

OpenStudy (anonymous):

This is what I did y=kx -2 = k + 7 k = -2/7 y = -2/7(-6) y = 1.7

OpenStudy (anonymous):

u will get x=21 . see... y=kx (suppose) so -2=k*7----------> k=-2/7 now if y=-6 then, -6=-2/7*x x=21

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

yeah enjoy! no problem :)

OpenStudy (cj49):

freaks sorry i missed that part.

OpenStudy (anonymous):

Thanks again guys.

OpenStudy (anonymous):

it's cool,it happens @cj49 and yeah pleasure to help @Yolo4mecuite